OpenStudy (anonymous):

A certain airplane has a speed of 283.6 km/h and is diving at an angle of θ = 29.0° below the horizontal when the pilot releases a radar decoy. The horizontal distance between the release point and the point where the decoy strikes the ground is d = 669 m. (a) How long is the decoy in the air? (b) How high was the release point?

OpenStudy (anonymous):

I got 16.2 seconds at a height of 2936.4, but I'm told I'm wrong.

OpenStudy (anonymous):

Did you include the y component of the airplane's speed in your calculation?

OpenStudy (anonymous):

I think so. To get time, I did: \[669divV _{x}\]

OpenStudy (anonymous):

to get height, I did 47.27m/s (sin29)+g(t^2)

OpenStudy (anonymous):

Just a minute, I'll draw something up.

OpenStudy (anonymous):

ooohhh. didn't divide by two for height.

OpenStudy (anonymous):

Ok, I have the initial x velocity at 78.8 m/s. That goes into this: \[s_x=vt\] or \[78.77\frac{m}{s}t = 669m\]. Now we have the air time, we should be able to get everything else.

OpenStudy (anonymous):

\[s_y = v_{yi}t + \frac{1}{2}at^2\]

OpenStudy (anonymous):

You shouldn't have to divide by two for the height, because you're only traveling one direction.

OpenStudy (anonymous):

8.49 seconds. How did you get your velocity? I just converted km/h to m/s.

OpenStudy (anonymous):

I got 744 m for height.

OpenStudy (anonymous):

I got 231m, having photo issues with my tablet, just a sec.

OpenStudy (anonymous):

\[78.77\left( \sin29 \right)+9.8m/s \left( 8.49^{2} \right)\]

OpenStudy (anonymous):

program is saying 231m and 8.41 seconds is wrong

OpenStudy (anonymous):

Aw crap. I just realized what I did wrong.

OpenStudy (anonymous):

8.49, i mean. Your drawing looks like mine.

OpenStudy (anonymous):

 I calculated 29 as the angle from the plane to the ground. It should be on top. The angle to calculate on is 61. That should fix it.

OpenStudy (anonymous):

Let me know if that works, gotta go. :(

OpenStudy (anonymous):

ahhhh!!!

OpenStudy (anonymous):

10 seconds and 897 feet.

OpenStudy (anonymous):

Cool. :)